# Multi-Frequency Progressive Refinement for Learned Inverse Scattering

### v1

@@ 1.1 Contributions & paper outline
    We consider alternative training strategies and find that the proposed progressive training strategy made possible by the MFISNet-Refinement architecture is important to its performance. (Section 5.5)
@@ 2 Problem Setup and Notation
The forward model for our imaging setup is implicitly defined by a PDE problem involving the Helmholtz equation. Let $x\in\mathbb{R}^{2}$ be the spatial variable. Suppose $u_{\mathrm{in}}(x;s)=e^{ikx\cdot s}$ is an incoming plane wave with direction $s\in\mathbb{S}^{1}$, wavelength $\lambda$, and angular wavenumber $k=2\pi/\lambda$. We normalize the problem’s units so this wave travels at speed $c_{0}\equiv 1$ in free space. The incoming wave interacts with a real-valued scattering potential $q(x)$ to produce an additive perturbation, called the scattered wave field $u_{\mathrm{scat}}[q](x;s)$. We define $q(x)=c_{0}^{2}/c^{2}(x)-1$ where $c(x)$ is the wave speed at $x$. The total wave field $u[q](x;s)=u_{\mathrm{scat}}[q](x;s)+u_{\mathrm{in}}(x;s)$ solves the following inhomogeneous Helmholtz equation, and the scattered wave field satisfies the Sommerfeld radiation boundary condition:
$\begin{cases}\Delta u[q](x;s)+k^{2}(1+q(x))u[q](x;s)=0&x\in\mathbb{R}^{2};\\[6.0pt] \dfrac{\partial u_{\mathrm{scat}}[q](x;s)}{\partial\|x\|}-iku_{\mathrm{scat}}[q](x;s)=o(\|x\|^{-1/2}),&\text{as $\|x\|\to\infty$.}\end{cases}$ (1)
We assume $q(x)$ is supported on a square domain $\Omega=[-0.5,0.5]^{2}$, and we work with $q\in\mathbb{R}^{N_{q}\times N_{q}}$, the discretization of $q(x)$ onto a regular grid with $(N_{q},N_{q})$ grid points. We place the receivers equally spaced around a large ring of radius $R\gg 1$, centered at the origin. We identify individual receivers by their unit-vector directions $r_{j}\in\mathbb{S}^{1}$. We compute the solution for a set of $N_{r}$ receiver directions and $N_{s}$ incoming wave directions $s_{\ell}$ equally-spaced about the unit circle. This results in a set of $(N_{r},N_{s})$ observations, $\{u_{\mathrm{scat}}[q](Rr_{j};s_{\ell})\}_{j,\ell\in[N_{r}]\times[N_{s}]}$, which we arrange in a data array $d_{k}\in\mathbb{C}^{N_{r}\times N_{s}}$. We call the mapping from $q$ to $d_{k}$ the forward model with incoming wave frequency $k$:
@@ 3.1 Background
The first result is that $\mathcal{F}_{k}$ becomes more nonlinear as the magnitude of the scatterer $\|q\|$ or the wavenumber $k$ increases. Indeed, the solution to (1) can be equivalently defined as the solution to the Lippmann-Schwinger integral equation:
$u_{\mathrm{scat}}[q](x;s)=k^{2}\int_{\Omega}G_{k}(\|x-x^{\prime}\|)q(x^{\prime})(u_{\mathrm{in}}(x^{\prime};s)+u_{\mathrm{scat}}[q](x^{\prime};s))\text{d}x^{\prime}$ (6)
@@
One way to view this nonlinearity is to interpret the Lippmann-Schwinger equation as a power series in $q(x)$ by iteratively substituting the value of $u_{\mathrm{scat}}[q](x)$ into its appearance on the right-hand side of (6). For example, performing this substitution once yields a linear and a quadratic term in $q(x)$ that are independent of $u_{\mathrm{scat}}[q]$, as well as a “remainder” term involving the unknown $u_{\mathrm{scat}}[q]$ that accounts for higher-order terms. This power series does not converge for general $q(x)$, but it helps illustrate which parts of the problem drive the nonlinearity of the operator $\mathcal{F}_{k}$: as $\|q\|$ or $k$ grow, the size of these nonlinear terms will also grow, and as a result $\mathcal{F}_{k}$ becomes highly nonlinear.
@@
We will refer to this linear approximation of the map from $q(x)$ to $d_{k}(r,s)$ as $F_{k}$. Note that $F_{k}q$ is proportional to the Fourier Transform of $q$ evaluated at frequency vectors of the form $k(r-s)$. Since $r,s\in\mathbb{S}^{1}$ range over the unit circle, the frequency vectors $k(r-s)$ take on values throughout a disk with radius $2k$ centered at the origin. Thus, evaluations of the linearized forward model $F_{k}q$ only contain the low-frequency components of $q$, while high-frequency components are in the kernel of $F_{k}$ (Chen, 1995).
@@
where $F_{k}^{*}$ is the adjoint of $F_{k}$. The operator $(F_{k}^{*}F_{k}+\mu I)^{-1}$ can be implemented as a two-dimensional spatial convolution (Khoo and Ying, 2019; Fan and Ying, 2022; Li et al., 2022), while novel network architectures have been proposed to emulate $F_{k}^{*}\in\mathbb{C}^{N_{q}^{2}\times N_{r}N_{s}}$ in a parameter-efficient manner. In particular, Fan and Ying (2022) and Zhang et al. (2023) suggest leveraging the rotational equivariance of the forward model to emulate $F_{k}^{*}$ with one-dimensional convolutions, after applying a far-field scaling and an appropriate coordinate transformation in (Fan and Ying, 2022, Equation 6)). We refer to the network described by Fan and Ying (2022) as FYNet.
@@ 3.2 Related Work
Several works in the wave scattering literature attempt to solve the inverse scattering problem by augmenting an optimization algorithm with components learned from data. At inference time, these methods require running an iterative optimization algorithm. Kamilov et al. (2017) develops a plug-and-play algorithm for inverse scattering, and show that various off-the-shelf denoisers can be applied as proximal operators. Zhao et al. (2023) use an encoder network paired with a network emulating the forward model and suggest optimizing a latent representation of the scattering potential using stochastic gradient descent. When only phaseless measurements of the scattered wave field $u_{\mathrm{scat}}$ are available, Deshmukh et al. (2022) propose a network unrolling proximal gradient descent, where the proximal operator is a neural network learned from data. For the inverse obstacle scattering problem in two dimensions, Zhou et al. (2023) propose using a fully-connected neural network to warm-start a Gauss-Newton algorithm. Ding et al. (2022) train a neural network to approximately invert a forward scattering process depending on temporal data, and use this approximate inverse as a nonlinear preconditioner for a nonlinear least squares optimization routine.
@@
Other methods propose to learn the inverse map directly from data. Recently, neural networks that are approximately invariant to discretization size have been proposed as methods of learning maps between general function spaces (Li et al., 2021b; Lu et al., 2021) and these general-purpose networks have been applied to inverse scattering (Ong et al., 2022). Other networks have been designed to invert the forward scattering model in particular; see Chen et al. (2020) for a broad review of such approaches. In our work, we are particularly interested in FBP-inspired methods. The work of Khoo and Ying (2019) leverages the approximate low-rank structure of scattering operators to design their SwitchNet network. Fan and Ying (2022) propose a data transformation that facilitates emulating the adjoint of the linearized forward model via 1D convolutions. In our work, we consider how to combine multiple such network blocks to invert the multi-frequency forward map in (3). One way of combining these blocks is to learn each adjoint operator $F_{k_{t}}^{*}$ as a separate neural network block and combine data to jointly emulate the learned filtering operators $(F^{*}_{k_{t}}F_{k_{t}}+\mu I)^{-1}$. This strategy is employed in Zhang et al. (2023), which is similar to our MFISNet-Parallel, but uses a different parameterization for the layers emulating $F_{k_{t}}^{*}$ and $(F^{*}_{k_{t}}F_{k_{t}}+\mu I)^{-1}$. Another strategy is to use the Wide-Band Butterfly Network (Li et al., 2022, 2021a), which hierarchically merges information from different frequencies in the network block emulating $F^{*}_{k_{t}}$. We provide more detail and commentary about methods of combining network blocks to form multi-frequency networks in Section 5.2.
@@ 4.1 Recursive Linearization
$\underset{\hat{q}}{\text{minimize~{}}}\lVert d_{k}-\mathcal{F}_{k}[\hat{q}]\rVert_{2}^{2}.$ (9)
@@ 5      $\hat{q}_{k_{t}}\leftarrow\hat{q}_{k_{t-1}}+\delta q_{k_{t}}$
where $\text{RefinementStep}_{k_{t}}(\hat{q}_{k_{t-1}},d_{k_{t}})$ refers to the update calculated for estimate $\hat{q}_{k_{t-1}}$ given data $d_{k_{t}}$ and can be implemented using a neural network. We will propose and discuss a network architecture in the next section.
@@ 4.2 Our Method
In (13), $\mathsf{LPF}_{2k_{t}}$ is a low-pass filter with approximate cutoff frequency $2k_{t}$, implemented as a Gaussian filter to avoid ringing artifacts; its frequency response is given by
@@ 4.3 Implementation of FYNet
We implement the inversion network described in (Fan and Ying, 2022) and call it FYNet. This method suggests applying the far-field scaling and a transformation from the receiver and source direction coordinates $(r,s)$ to a new set of variables as summarized in (Fan and Ying, 2022, Equation (6)), which we perform using bicubic interpolation. We split the complex-valued input into real and imaginary parts along a channel dimension. To implement the action of the adjoint operator $F_{k}^{*}$, we use a composition of three 1-dimensional convolutional layers. We implement the convolutional layers as learnable in the Fourier domain because the expression for $F_{k}^{*}$ derived in (Fan and Ying, 2022) is local in frequency, but not space. To implement the filtering operator $\left(F_{k}^{*}F_{k}+\mu I\right)^{-1}$, we use a composition of three 2-dimensional convolutional layers. All layers but the last one use ReLU activations. This network outputs an estimate of the scattering potential on a regular polar grid, in coordinates $(\rho,\phi)$. Following (Fan and Ying, 2022), we train the network by minimizing the difference between predictions and targets on the polar grid. We transform to Cartesian coordinates for visualization and computing final test statistics.
@@ 5.1 Dataset and Data Generation
Figure 3: Figure 3(a) shows a typical example from our distribution of scattering potentials, drawn from the test set. Our distribution of scattering potentials has a random low-frequency background field, occluded by piecewise constant geometric shapes. Figures 3(b) and 3(c) show the output of the forward model applied to this scattering potential. In these plots, incident wave frequencies $4\pi$ and $32\pi$ were used, respectively. The real part of $u_{\mathrm{scat}}$ is shown.
@@
To implement the forward model, we implement a numerical PDE solver to compute solutions of (1). We implement this by discretizing the scattering domain $\Omega$ with a $(N_{q},N_{q})$ regular grid, with $N_{q}=192$. We transform (1) into the Lippmann-Schwinger integral equation and recast the latter as a sparse linear system, which we solve with an iterative method to approximate convergence. This formulation allows us to compute the solution $u_{\mathrm{scat}}$ on a large, distant ring placed at radius $R=100$. We compute the solution at $N_{r}=192$ equally-spaced positions on this ring, and we repeat this process for $N_{s}=192$ equally-spaced source directions. We use a Pytorch implementation of GMRES (Saad and Schultz, 1986) through the CoLA library (Potapczynski et al., 2023) to facilitate batched computation over multiple source directions, fast sparse linear system solves, and GPU acceleration; altogether, evaluating the full forward model for a given $q$ only takes about 5 seconds using our implementation, using one $\text{NVIDIA}^{\text{\textregistered}}$ A40 GPU. Figures 3(b) and 3(c) show the solution $u_{\mathrm{scat}}$ produced by our implementation for high and low incident wave frequencies.
We use bicubic interpolation to implement the data transformation described in (Fan and Ying, 2022, Equation (6)), resulting in a transformed grid of size $(N_{m},N_{h})=(192,96)$ for the input data. The FYNet blocks reconstruct images on a regular polar grid with $(N_{\rho},N_{\phi})=(96,192)$ pixels, and the radial dimension of our polar grid extends to $\rho_{\max}=0.5$. Finally, we use bicubic interpolation to transform the model’s outputs to the $(N_{q},N_{q})$ Cartesian grid for final visualization and error measurement.

### v2

@@ 1.1 Contributions & paper outline
    We consider alternative training strategies and find that MFISNet-Refinement is robust to the choice of training method, suggesting the majority of improvement is due to the architecture. (Section 5.5)
  5. 5.
@@ 2 Problem Setup and Notation
Figure 1: Geometry of the inverse scattering problem. An incident plane wave $u_{\mathrm{in}}(x;s_{\ell})$ coming from source direction $s_{\ell}$ interacts with the scattering potential $q(x)$. The resulting scattered wave field $u_{\mathrm{scat}}[q](x;s_{\ell})$ is recorded by a receiver $r_{j}$.
The forward model for our imaging setup is implicitly defined by a PDE problem involving the Helmholtz equation. See Figure 1 for a diagram of the geometry of the problem. Let $x\in\mathbb{R}^{2}$ be the spatial variable. Suppose $u_{\mathrm{in}}(x;s)=e^{ikx\cdot s}$ is an incoming plane wave with direction $s\in\mathbb{S}^{1}$, wavelength $\lambda$, and angular wavenumber $k=2\pi/\lambda$. We normalize the problem’s units so this wave travels at speed $c_{0}\equiv 1$ in free space. The incoming wave interacts with a real-valued scattering potential $q(x)$ to produce an additive perturbation, called the scattered wave field $u_{\mathrm{scat}}[q](x;s)$. We define $q(x)=c_{0}^{2}/c^{2}(x)-1$ where $c(x)$ is the wave speed at $x$. The total wave field $u[q](x;s)=u_{\mathrm{scat}}[q](x;s)+u_{\mathrm{in}}(x;s)$ solves the following inhomogeneous Helmholtz equation, and the scattered wave field satisfies the Sommerfeld radiation boundary condition:
$\begin{cases}\Delta u[q](x;s)+k^{2}(1+q(x))u[q](x;s)=0&x\in\mathbb{R}^{2};\\[6.0pt] \dfrac{\partial u_{\mathrm{scat}}[q](x;s)}{\partial\|x\|_{2}}-iku_{\mathrm{scat}}[q](x;s)=o(\|x\|_{2}^{-1/2}),&\text{as $\|x\|_{2}\to\infty$.}\end{cases}$ (1)
We assume $q(x)$ is supported on a square domain $\Omega=[-0.5,0.5]^{2}$, and we work with $q\in\mathbb{R}^{N_{q}\times N_{q}}$, the discretization of $q(x)$ onto a regular grid with $(N_{q},N_{q})$ grid points. We place the receivers equally spaced around a large ring of radius $R\gg 1$, centered at the origin. We identify individual receivers by their unit-vector directions $r_{j}\in\mathbb{S}^{1}$. We compute the solution for the same set of $N_{r}$ receiver points and $N_{s}$ incoming source directions, where $N_{r}=N_{s}$ and the grid points are equally distributed about the unit circle. This results in a set of $(N_{r},N_{s})$ observations, $\{u_{\mathrm{scat}}[q](Rr_{j};s_{\ell})\}_{j,\ell\in[N_{r}]\times[N_{s}]}$, which we arrange in a data array $d_{k}\in\mathbb{C}^{N_{r}\times N_{s}}$. We call the mapping from $q$ to $d_{k}$ the forward model with incoming wave frequency $k$:
@@ 3.1 Background
The first result is that $\mathcal{F}_{k}$ becomes more nonlinear as the magnitude of the scatterer $\|q\|_{2}$ or the wavenumber $k$ increases. Indeed, the solution to (1) can be equivalently defined as the solution to the Lippmann-Schwinger integral equation:
$u_{\mathrm{scat}}[q](x;s)=k^{2}\int_{\Omega}G_{k}(\|x-x^{\prime}\|_{2})q(x^{\prime})(u_{\mathrm{in}}(x^{\prime};s)+u_{\mathrm{scat}}[q](x^{\prime};s))\text{d}x^{\prime}$ (6)
@@
One way to view this nonlinearity is to interpret the Lippmann-Schwinger equation as a power series in $q(x)$ by iteratively substituting the value of $u_{\mathrm{scat}}[q](x)$ into its appearance on the right-hand side of (6). For example, performing this substitution once yields a linear and a quadratic term in $q(x)$ that are independent of $u_{\mathrm{scat}}[q]$, as well as a “remainder” term involving the unknown $u_{\mathrm{scat}}[q]$ that accounts for higher-order terms:
$u_{\text{scat}}[q](x;s)$ $=k^{2}\int_{\Omega}G_{k}(\|x-x^{\prime}\|_{2})q(x^{\prime})u_{\text{in}}(x^{\prime};s)\text{d}x^{\prime}$
$+$ $k^{4}\int_{\Omega}G_{k}(\|x-x^{\prime}\|_{2})q(x^{\prime})\int_{\Omega}G_{k}(\|x^{\prime}-x^{\prime\prime}\|_{2})q(x^{\prime\prime})u_{\text{in}}(x^{\prime\prime};s)\text{d}x^{\prime\prime}\text{d}x^{\prime}$
[…]
This power series does not converge for general $q(x)$, but it helps illustrate which parts of the problem drive the nonlinearity of the operator $\mathcal{F}_{k}$: as $\|q\|_{2}$ or $k$ grow, the size of these nonlinear terms will also grow, and as a result $\mathcal{F}_{k}$ becomes increasingly nonlinear.
@@
We will refer to this linear approximation of the map from $q(x)$ to $d_{k}(r,s)$ as $F_{k}$. Note that $F_{k}q$ is proportional to the Fourier Transform of $q$ evaluated at frequency vectors of the form $k(r-s)$. Since $r,s\in\mathbb{S}^{1}$ range over the unit circle, the frequency vectors $k(r-s)$ take on values throughout a disk with radius $2k$ centered at the origin. Thus, evaluations of the linearized forward model, $F_{k}q$, only contain the low-frequency components of $q$, while high-frequency components of $q$ are in the kernel of the linearized forward model $F_{k}$ (Chen, 1995).
@@
where $F_{k}^{*}$ is the adjoint of $F_{k}$ and $\mu$ is a regularization parameter that stabilizes the inversion of $F_{k}^{*}F_{k}$. The operator $(F_{k}^{*}F_{k}+\mu I)^{-1}$ can be implemented as a two-dimensional spatial convolution (Khoo and Ying, 2019; Fan and Ying, 2022; Li et al., 2022), while novel network architectures have been proposed to emulate $F_{k}^{*}\in\mathbb{C}^{N_{q}^{2}\times N_{r}N_{s}}$ in a parameter-efficient manner. In particular, Fan and Ying (2022) and Zhang et al. (2023) suggest leveraging the rotational equivariance of the forward model to emulate $F_{k}^{*}$ with one-dimensional convolutions, after applying a far-field scaling and an appropriate coordinate transformation in (Fan and Ying, 2022, Equation 6)). We refer to the network described by Fan and Ying (2022) as FYNet.
@@ 3.2 Related Work
Several works in the wave scattering literature attempt to solve the inverse scattering problem by augmenting an optimization algorithm with components learned from data. At inference time, these methods require running an iterative optimization algorithm. Kamilov et al. (2017) develops a plug-and-play algorithm for inverse scattering, and show that various off-the-shelf denoisers can be applied as proximal operators. Zhao et al. (2023) use an encoder network paired with a network emulating the forward model and suggest optimizing a latent representation of the scattering potential using stochastic gradient descent. When only phaseless measurements of the scattered wave field $u_{\mathrm{scat}}$ are available, Deshmukh et al. (2022) propose a network unrolling proximal gradient descent, where the proximal operator is a neural network learned from data. For the inverse obstacle scattering problem in two dimensions, Zhou et al. (2023) propose using a fully-connected neural network to warm-start a Gauss-Newton algorithm. Ding et al. (2022) train a neural network to approximately invert a forward scattering process depending on temporal data, and use this approximate inverse as a nonlinear preconditioner for a nonlinear least squares optimization routine. In concurrent work, Zhang et al. (2024) use diffusion sampling to reconstruct scattering potentials and quantify uncertainty of the reconstructions.
@@
Other methods propose to learn the inverse map directly from data. Recently, neural networks that are approximately invariant to discretization size have been proposed as methods of learning maps between general function spaces (Li et al., 2021b; Lu et al., 2021) and these general-purpose networks have been applied to inverse scattering (Ong et al., 2022). Other networks have been designed to invert the forward scattering model in particular; see Chen et al. (2020) for a broad review of such approaches. In our work, we are particularly interested in FBP-inspired methods. The work of Khoo and Ying (2019) leverages the approximate low-rank structure of scattering operators to design their SwitchNet network. Fan and Ying (2022) propose a data transformation that facilitates emulating the adjoint of the linearized forward model via 1D convolutions. In our work, we consider how to combine multiple such network blocks to invert the multi-frequency forward map in (3). One way of combining these blocks is to learn each adjoint operator $F_{k}^{*}$ as a separate neural network block and combine data to jointly emulate the learned filtering operators $(F^{*}_{k}F_{k}+\mu I)^{-1}$. This strategy is employed in Zhang et al. (2023), which is similar to our MFISNet-Parallel, but uses a different parameterization for the layers emulating $F_{k}^{*}$ and $(F^{*}_{k}F_{k}+\mu I)^{-1}$. Another strategy is to use the Wide-Band Butterfly Network (Li et al., 2022, 2021a), which hierarchically merges information from different frequencies in the network block emulating $F^{*}_{k}$. We provide more detail and commentary about methods of combining network blocks to form multi-frequency networks in Section 5.2.
@@ 4.1 Recursive Linearization
$\underset{\hat{q}}{\text{argmin~{}}}\lVert d_{k}-\mathcal{F}_{k}[\hat{q}]\rVert_{2}^{2}.$ (10)
@@ 5      $\hat{q}_{k_{t}}\leftarrow\hat{q}_{k_{t-1}}+\delta q_{k_{t}}$
where $\text{RefinementStep}_{k_{t}}(\hat{q}_{k_{t-1}},d_{k_{t}})$ refers to the update calculated for estimate $\hat{q}_{k_{t}}$ given data $d_{k_{t}}$ and can be implemented using a neural network. We will propose and discuss a network architecture in the next section.
@@ 4.2 Our Method
In (14), $\mathsf{LPF}_{2k_{t}}$ is a low-pass filter with approximate cutoff frequency $2k_{t}$, implemented as a Gaussian filter to avoid ringing artifacts; its frequency response is given in the Fourier domain by
@@ 4.3 Implementation of FYNet
We implement the inversion network described in (Fan and Ying, 2022) and call it FYNet. This method suggests applying the far-field scaling and a transformation from the receiver and source direction coordinates $(r,s)$ to a new set of variables as summarized in (Fan and Ying, 2022, Equation (6)), which we perform using bicubic interpolation. We describe this transformation in Equation 16. We split the complex-valued input into real and imaginary parts along a channel dimension. To implement the action of the adjoint operator $F_{k}^{*}$, we use a composition of three 1-dimensional convolutional layers. We implement the convolutional layers as learnable in the Fourier domain because the expression for $F_{k}^{*}$ derived in (Fan and Ying, 2022) is local in frequency, but not space. To implement the filtering operator $\left(F_{k}^{*}F_{k}+\mu I\right)^{-1}$, we use a composition of three 2-dimensional convolutional layers. All layers but the last one use ReLU activations. This network outputs an estimate of the scattering potential on a regular polar grid, in coordinates $(\rho,\phi)$. Following (Fan and Ying, 2022), we train the network by minimizing the difference between predictions and targets on the polar grid. We transform to Cartesian coordinates for visualization and computing final test statistics.
@@ 5.1 Dataset and Data Generation
Figure 4: Figure 4(a) shows a typical example from our distribution of scattering potentials, drawn from the test set. Our distribution of scattering potentials has a random low-frequency background field, occluded by piecewise constant geometric shapes. The bottom-left corner shows the wavelength of two incident waves with frequencies $k=4\pi$ and $k=32\pi$. Figures 4(b) and 4(c) show the output of the forward model applied to this scattering potential, using these incident frequencies. The real part of $u_{\mathrm{scat}}$ is shown.
@@
To implement the forward model, we implement a numerical PDE solver to compute solutions of (1). We implement this by discretizing the scattering domain $\Omega$ with a $(N_{q},N_{q})$ regular grid, with $N_{q}=192$. We transform (1) into the Lippmann-Schwinger integral equation and recast the latter as a sparse linear system, which we accelerate using fast Fourier transforms and hardware acceleration. We solve this linear system with GMRES (Saad and Schultz, 1986) implemented by SciPy (Virtanen et al., 2020) to a relative tolerance of $10^{-2}$. This formulation allows us to compute the solution $u_{\mathrm{scat}}$ on a large, distant ring placed at radius $R=100$. We compute the solution at $N_{r}=192$ equally-spaced positions on this ring, and we repeat this process for each of the $N_{s}=192$ equally-spaced source directions. The sources and receivers are located on the same grid. The runtime for evaluating the full forward model for a given $q$ at the lowest and highest incident wave frequencies considered requires 5 and 220 seconds respectively, using one $\text{NVIDIA}^{\text{\textregistered}}$ A40 GPU. Figures 4(b) and 4(c) show the solution $u_{\mathrm{scat}}$ produced by our implementation for low and high incident wave frequencies, respectively.
Fan and Ying (2022) use a coordinate transformation of the far-field scattering data from coordinates $(r,s)$ to $(m,h)$, where
$m:=\frac{r+s}{2}\quad\text{and}\quad h:=\frac{r-s}{2}$ (16)
as described in (Fan and Ying, 2022, Equation (6)). The variable $m$ ranges from 00 to $2\pi$, while $h$ ranges from $-\pi/2$ to $\pi/2$. As in Fan and Ying (2022), we choose $N_{m}=192$ and $N_{h}=96$ to keep the angular sampling frequency fixed for all angular variables considered in the problem. We perform this transformation using bicubic interpolation, and the resulting grid has dimensions $(N_{m},N_{h})=(192,96)$.
The FYNet blocks reconstruct images on a regular polar grid with $(N_{\rho},N_{\phi})=(96,192)$ pixels, and the radial dimension of our polar grid extends to $\rho_{\max}=0.5$. Finally, we use bicubic interpolation to transform the model’s outputs to the $(N_{q},N_{q})$ Cartesian grid for final visualization and error measurement.